OpenStudy (anonymous):

In parallelogram DEFG where they meet in the middle at H, DH=x+1, HF=3y, GH=3x-4, and HE=5y+1. Find the values of x and y

OpenStudy (anonymous):

please explain also

OpenStudy (anonymous):

where is the link there is none copy paste it

OpenStudy (anonymous):

It won't let me copy and paste it

OpenStudy (anonymous):

I think that x=5 and y=2 but i'm not sure

OpenStudy (anonymous):

If I understand it correctly that means DH=GH and HE=HF so setup the equations x+1=3x-4 3y=5y+1 then solve for each. I got X=2.5 and Y=-0.5

OpenStudy (anonymous):

according to the picture of the parallelogram DH=HF and GH=HE :/

OpenStudy (anonymous):

oh nm I did it wrong

OpenStudy (anonymous):

you're right @EJCeballos thank you

OpenStudy (anonymous):

Upper left being D, upper right being E, bottom right being G ,and bottom left being F?